1.     Briefly explain any THREE (3) different approaches that are used to measure Gross Domestic Product (GDP). (6 MARK)

2.     Distinguish between ‘Gross Domestic Product’ and ‘Gross National Product’.( 4marks)

3.    The table below shows the value of nominal GDP and price index of a country:

Year price index Nominal GDP

2008 100 RM 40,000 million

2009 107 RM 48,000 million

(i)              Calculate the value of Real GDP in 2008 and 2009. (6marks)

(ii)              Calculate the rate of economic growth between 2008and 2009. (3marks)


1
Expert's answer
2021-03-22T08:07:08-0400
"solution"

A] expenditure approach: its the total spending on all final goods and services (Consumption goods and services (C) + Gross Investments (I) + Government Purchases (G) + (Exports (X) – Imports (M)) GDP = C + I + G + (X-M).

income approach: GDP based on the income approach is calculated by adding up the factor incomes to the factors of production in the society.

income="R+W+Ir+Pr"

output approach: GDP is calculated using the output approach by summing the value of sales of goods and adjusting (subtracting) for the purchase of intermediate goods to produce the goods sold.


B] Gross Domestic Product (GDP) is used to calculate the total economic value of all the final goods and services produced within a country’s borders within a specific period of time usually one year , while Gross National Product (GNP) is used to calculate the total economic value of the goods and services produced by the residents of a country, no matter their location thus includes their foreign assets too.


C] (i) the nominal GDP of year 2008 is given as "40,000 million" :

let 2008 be the base year therefore the real GDP will be assumed to be equal to the nominal GDP

therefore it is "\$40,000 \ million"


the real GDP of 2009


"\frac{48000}{107}=448.598=448.60\\"


"448.60\times100=\$44,860 \ million"


(ii)the rate of economic growth is


GDP Def="[\frac{nominal \ GDP}{real\ GDP}] \times 100"


"\frac{48000}{44860}\times100\%=106.99=107"


its "107-100=7\%"